OpenStudy (anonymous):

Use the substitution method to solve the following system of linear equations: x + y = 6 x = –y + 2 A. The solution is (4, 2). B. There are an infinite number of solutions. C. There is no solution. D. The solution is (2, 4).

OpenStudy (btaylor):

Since you know what x is (in terms of y), plug that into the given value of x. So:\[x+y=6 \rightarrow (-y+2)+y=6 \rightarrow y-y+2=6 \rightarrow 2=6 \rightarrow ???\]Since we get an untrue statement (2=6), there is no solution, which means that no combination of x and y will result in an equality. Basically if you were to graph the two lines defined by the equation, they are parallel, so they never intersect (at a solution)